# CaSTFormer: Causal Spatio‐Temporal Transformer for Driving Intention Prediction

### v1

@@ Introduction
Brain4Cars: The Brain4Cars dataset (Jain et al. 2016a) comprises exterior (480 $\times$ 720) and interior (1088 $\times$ 1920) videos of up to 5-second segments, refined to 594 valid events after excluding incomplete or unsynchronized samples. Each video is uniformly sampled to 150 frames, extracting the 5-second segment preceding the maneuver. Interior frames are cropped to 900 $\times$ 800, resized to 112 $\times$ 112, and encoded by a MobileFaceNet yielding 64-D features. Exterior frames undergo RAFT optical flow computation, are resized to 144 $\times$ 96, and then processed by ResNet-18 to produce 32-D features. Appending a smoothed speed signal yields 65-D (internal) and 33-D (external) vectors. These vectors are linearly projected, positionally encoded, and passed through a Transformer encoder to obtain temporal representations for CaSTFormer. The dataset spans highway and urban settings with five maneuver classes: straight, left turn, right turn, left lane change, and right lane change.
Implementation Details
Our proposed CaSTFormer was implemented by PyTorch and experiments were performed on a server with six NVIDIA RTX 2080 Ti GPUs. The model was trained end‐to‐end on Brain4Cars using the Adam optimizer (initial learning rate $1\times 10^{-3}$) for 160 epochs with a batch size of 16. During training, each input comprised a chunk of frames randomly sampled from the 5-second pre-maneuver segment; at inference, chunks were obtained via uniform sampling. The weight $\alpha$ in the unified loss was empirically set to 0.1. Model performance was evaluated using 5-fold cross-validation.
Evaluation Protocols
In driving intention prediction, straight driving is considered background, and only turns and lane changes are treated as target events. To evaluate our CaSTFormer model, we define the following metrics based on the predictions: true positives (TP: correctly predicted maneuvers), false positives (FP: maneuvers predicted incorrectly as another maneuver), false positive predictions (FPP: predicting a maneuver when none occurred), and missing predictions (MP: failing to detect an actual maneuver). Given the set of all behaviors $\mathcal{G}$ and target maneuvers $\mathcal{G}^{\prime}=\mathcal{G}\setminus\{\mathrm{straight}\}$, Precision (Pr), Recall (Re), and F1-score are computed as follows:
$\mathrm{Pr}$ $=\frac{1}{|\mathcal{G}^{\prime}|}\sum_{m\in\mathcal{G}^{\prime}}\frac{TP_{m}}{TP_{m}+FP_{m}+FPP_{m}},$ (13)
$\mathrm{Re}$ $=\frac{1}{|\mathcal{G}^{\prime}|}\sum_{m\in\mathcal{G}^{\prime}}\frac{TP_{m}}{TP_{m}+MP_{m}},\quad F_{1}=\frac{2*\mathrm{Pr}*\mathrm{Re}}{\mathrm{Pr}+\mathrm{Re}}.$
Comparison with State-of-the-art Methods
Table 2: Comparison of our CaSTFormer against other end-to-end methods on the Brain4Cars dataset, using internal and external streams, with F1-score (%) and parameters (M).
Method Inside Outside F1-score (%) Param. (M)
@@
Rong (Rong, Akata, and Kasneci 2020) $\checkmark$ 75.5 46.2+162
@@
CaSTFormer (Ours) $\checkmark$ $\checkmark$ 98.6 14.53+5.3
Table 1 provides a systematic comparison of both single- and multi-modal methods on the Brain4Cars dataset. Notably, our camera-only CaSTFormer variant achieves an F1-score of 97.6% (precision 96.7%, recall 98.5%), markedly surpassing all previous single-modality methods such as DCNN (92.1%) and CF-LSTM (92.1%). When enriched with speed information, CaSTFormer attains a new state-of-the-art F1-score of 98.6% (precision 98.7%, recall 98.5%), outperforming the best prior multi-modal model, FedPRM (95.2% F1), by 3.4%. These results demonstrate that CaSTFormer not only establishes a new performance standard but does so with fewer sensor inputs, highlighting its efficiency and robustness in driving intention prediction. Figure 4 presents the confusion matrices of our CaSTFormer and TIFN (Guo et al. 2023). CaSTFormer yields a sharper diagonal and substantially fewer off-diagonal entries, demonstrating its superior discrimination of similar maneuvers and reduced false predictions.
Table 3: F1-scores on the Brain4Cars dataset for evaluation on video segments truncated 1–4 s prior to action onset.
@@
Rong (Rong, Akata, and Kasneci 2020) (in) 75.7 73.1 68.6 58.5 48.2
Rong (Rong, Akata, and Kasneci 2020) (out) 66.4 62.4 47.0 38.8 38.9
Rong (Rong, Akata, and Kasneci 2020) (both) 84.3 78.9 70.6 60.3 53.4
@@
CaSTFormer (Ours) 98.6 97.4 90.1 78.4 63.7
Detailed comparative results between our method and other end-to-end approaches on full-video inputs appear in Table 2. As the optical-flow algorithm lies outside the core prediction pipeline, its parameters are listed separately. Our model achieves superior recognition performance with only a marginal increase in model size, demonstrating its compact efficiency.
In addition to evaluating F1-scores on complete videos (–5 s to 0 s, where 0 s marks driver action), we assessed early-warning capability by truncating observation windows at –1 s, –2 s, –3 s, and –4 s. As shown in Table 3, prediction accuracy declines nearly linearly with shorter observations, highlighting increased uncertainty at longer forecast horizons. This result reflects the intrinsic trade-off between early intervention and predictive accuracy. Our CaSTFormer consistently achieves superior performance across all truncated settings, demonstrating its robustness in driving intention prediction.
@@
Ablation Study
Effect of components in CaSTFormer.
To evaluate the contribution of each component in CaSTFormer, we conducted systematic ablation studies with results presented in Table 4. Using a dual-stream Transformer baseline (T-Base) that independently encodes internal and external features through simple concatenation, we progressively introduced the RSF, CPE and FSN modules to assess their individual impact on F1-score performance. We further explored various combined configurations to understand module interactions and their cumulative effects. The experiments reveal consistent performance improvements across all configurations, validating that each module makes meaningful contributions to driving intention prediction while their synergistic combination yields additive enhancement benefits.
Table 4: F1-scores on the Brain4Cars dataset between the dual-stream Transformer baseline and progressively augmented variants that introduce RSF, CPE and FSN.
@@
T-Base 95.8 94.2 85.4 73.7 63.2
T-Base+RSF 97.1 95.6 87.1 75.2 61.9
T-Base+CPE 97.0 95.4 86.9 74.9 61.1
T-Base+FSN 96.6 94.9 86.3 73.9 62.6
T-Base+RSF+CPE 97.4 95.7 87.4 75.9 60.3
T-Base+RSF+FSN 98.0 96.7 88.5 76.8 65.6
T-Base+CPE+FSN 97.8 96.4 88.0 76.2 62.4
T-Base+RSF+CPE +FSN(CaSTFormer) 98.6 97.4 90.1 78.4 63.7
Figure 5: F1‐scores (%) of different attention designs on the Brain4Cars dataset.
Effect of attention design in CaSTFormer.
To evaluate the effect of different attention mechanisms on driving intention prediction, we conducted ablation experiments on the Brain4Cars dataset comparing three attention schemes, as summarized in Table 5. The results indicate that our proposed Bidirectional Dependency Attention (BDA) more effectively captures spatio-temporal correlations between interior and exterior streams while suppressing noise, thereby demonstrating its superiority in isolating dynamic cues and enhancing overall model robustness.
Table 5: F1‐scores (%) of different the sequential correspondence orders in CaSTFormer on the Brain4Cars dataset.
Module Sequence F1‐score (%)
[-5,0] [-5,-1] [-5,-2] [-5,-3] [-5,-4]
@@
[…]
Effect of module sequencing in CaSTFormer.
[…]
@@
To assess the influence of the intention‑loss weight $\alpha$ in our composite objective, we conducted an ablation study by varying $\alpha$ to identify the optimal balance between core spatio‑temporal feature learning and dedicated intention supervision. The results in Fig. 6 show that setting $\alpha=0.10$ produces the highest F1‑scores across most anticipation horizons, because this value injects enough gradient from the intention loss to enrich the discriminative power of spatio‑temporal embeddings with causal cues while preserving the stability and generalization capacity of the primary classification objective. Smaller values of $\alpha$ under‑utilize intention supervision and fail to capture these pre‑action signals, whereas larger values allow the intention loss to dominate optimization, creating gradient conflicts that degrade both spatio‑temporal coherence and overall anticipation accuracy.
@@ Conclusion
In this paper, we propose CaSTFormer, an interpretable prototype-driven causal spatio-temporal transformer for driving intention prediction. Our approach extracts dual‐stream interior and exterior features and processes them through a structured pipeline to align and fuse multi‐modal representations. Extensive experiments on the Brain4Cars dataset demonstrate that CaSTFormer achieves state‐of‐the‐art prediction accuracy with only a marginal increase in model size, validating its efficiency and suitability for real‐time driver assistance.

### v2

@@ Introduction
Brain4Cars: The Brain4Cars dataset (Jain et al. 2016a) comprises exterior (480 $\times$ 720) and interior (1088 $\times$ 1920) videos of up to 5-second segments, refined to 594 valid events after excluding incomplete or unsynchronized samples. Each video is uniformly sampled to 150 frames, extracting the 5-second segment preceding the maneuver. Interior frames are cropped to 900 $\times$ 800, resized to 112 $\times$ 112, and encoded by a MobileFaceNet yielding 64-D features. Exterior frames undergo Farneback optical flow computation, are resized to 144 $\times$ 96, and then processed by ResNet-18 to produce 32-D features. Appending a smoothed speed signal yields 65-D (interior) and 33-D (exterior) vectors. These vectors are linearly projected, positionally encoded, and passed through a Transformer encoder to obtain temporal representations for CaTFormer. The dataset spans highway and urban settings with five maneuver classes: straight, left turn, right turn, left lane change, and right lane change.
Method In Out F1 (%) Param. (M)
@@
Rong (Rong et al. 2020) $\checkmark$ 75.5 46.2+162
@@
CaTFormer (Ours) $\checkmark$ $\checkmark$ 98.6 14.53+5.3
Table 2: Comparison of our CaTFormer against other end-to-end methods on the Brain4Cars dataset, using interior and exterior streams, with F1-score (%) and parameters (M).
Implementation Details
Our proposed CaTFormer was implemented by PyTorch, and experiments were performed on a server with six NVIDIA RTX 2080 Ti GPUs. The model was trained end‐to‐end on Brain4Cars using the Adam optimizer (initial learning rate $1\times 10^{-3}$) for 160 epochs with a batch size of 16. During training, each input comprised a chunk of frames randomly sampled from the 5-second pre-maneuver segment. When testing, chunks were obtained via uniform sampling. The weight $\alpha$ in the unified loss was empirically set to 0.1. Model performance was evaluated using 5-fold cross-validation.
Evaluation Protocols
In driving intention prediction, straight driving is considered background, and only turns and lane changes are treated as target events. To evaluate our CaTFormer model, we define the following prediction-based metrics: true positives (TP: correctly predicted maneuvers), false positives (FP: maneuvers misclassified as another maneuver), false optimistic predictions (FPP: predicting a maneuver when none occurred), and missing predictions (MP: failing to detect an actual maneuver). Given the set of all behaviors $\mathcal{G}$ and target maneuvers $\mathcal{G}^{\prime}=\mathcal{G}\setminus\{\mathrm{straight}\}$, Precision (Pr), Recall (Re), and F1-score are computed as follows:
$\mathrm{Pr}$ $=\frac{1}{|\mathcal{G}^{\prime}|}\sum_{m\in\mathcal{G}}\frac{TP_{m}}{TP_{m}+FP_{m}+FPP_{m}},$ (13)
$\mathrm{Re}$ $=\frac{1}{|\mathcal{G}^{\prime}|}\sum_{m\in\mathcal{G}}\frac{TP_{m}}{TP_{m}+MP_{m}},\ F_{1}=\frac{2*\mathrm{Pr}*\mathrm{Re}}{\mathrm{Pr}+\mathrm{Re}}.$
@@
Rong (Rong et al. 2020) (in) 75.7 73.1 68.6 58.5 48.2
Rong (Rong et al. 2020) (out) 66.4 62.4 47.0 38.8 38.9
Rong (Rong et al. 2020) (both) 84.3 78.9 70.6 60.3 53.4
@@
CaTFormer (Ours) 98.6 97.4 90.1 78.4 63.7
Table 3: F1-scores on the Brain4Cars dataset for evaluation on video segments truncated 1–4 s before action onset.
Comparison with State-of-the-art Methods
Table 1 provides a systematic comparison of both single- and multi-modal methods on the Brain4Cars dataset. Notably, our camera-only CaTFormer variant achieves an F1-score of 97.6% (precision 96.7%, recall 98.5%), markedly surpassing all previous single-modality methods such as DCNN (92.1%) and CF-LSTM (92.1%). When enriched with speed information, CaTFormer attains a new state-of-the-art F1-score of 98.6% (precision 98.7%, recall 98.5%), outperforming the best prior multi-modal model, FedPRM (95.2% F1), by 3.4%. These results demonstrate that CaTFormer not only establishes a new performance standard but does so with fewer sensor inputs, highlighting its efficiency and robustness in driving intention prediction. Fig. 4 presents the confusion matrices of our CaTFormer and TIFN (Guo et al. 2023). CaTFormer yields a sharper diagonal and substantially fewer off-diagonal entries, demonstrating its superior discrimination of similar maneuvers and reduced false predictions. Detailed comparative results between our method and other end-to-end approaches on full-video inputs appear in Table 2. As the optical-flow algorithm lies outside the core prediction pipeline, its parameters are listed separately. Our model achieves superior recognition performance with only a marginal increase in model size, demonstrating its compact efficiency.
In addition to evaluating F1-scores on complete videos (–5 s to 0 s, where 0 s marks driver action), we assessed early-warning capability by truncating observation windows at –1 s, –2 s, –3 s, and –4 s. As shown in Table 3, prediction accuracy declines nearly linearly with shorter observations, highlighting increased uncertainty at longer forecast horizons. This result reflects the intrinsic trade-off between early intervention and predictive accuracy. Our CaTFormer consistently achieves superior performance across all truncated settings, demonstrating its robustness in driving intention prediction.
@@
Base 95.8 94.2 85.4 73.7 63.2
Base+R 97.1 95.6 87.1 75.2 61.9
Base+C 97.0 95.4 86.9 74.9 61.1
Base+F 96.6 94.9 86.3 73.9 62.6
Base+R+C 97.4 95.7 87.4 75.9 60.3
Base+R+F 98.0 96.7 88.5 76.8 65.6
Base+C+F 97.8 96.4 88.0 76.2 62.4
CaTFormer (R+C+F) 98.6 97.4 90.1 78.4 63.7
Table 4: F1‐scores on the Brain4Cars dataset for the dual-stream Transformer baseline (Base) and its variants augmented with RDF (R), CRE (C), and FSN (F).
@@
Result Visualization
Fig. 5 demonstrates that the model employs a temporal attention mechanism to realize a full reasoning path from dynamic event understanding to static decision attribution. Temporally, (a) and (b) delineate broad, task‑relevant event windows (e.g., the lane‑change interval), whereas (c) and (d) concentrate on a small set of decisive frames, accentuating discriminative cues and pinpointing instantaneous triggers. This sequence closely mirrors human cognition in which one perceives an event in its entirety before pinpointing its core cause. In the spatial dimension, (e) and (f) anchor the model’s reasoning in semantically relevant regions, including the driver’s facial state inside the cabin and the key road environment outside, entirely consistent with real‑world logic. This process vividly demonstrates how the model integrates critical cues to arrive at a reliable judgment, thereby substantiating the soundness of its decision‑making.
Ablation Study
Effect of components in CaTFormer.
To evaluate the contribution of each component in CaTFormer, we conduct systematic ablation studies, as shown in Table 4. Starting from a dual-stream Transformer baseline (Base) that performs late fusion via feature concatenation, we progressively add RDF, CRE, and FSN to measure their impact on F1-score. We further explored various combinations to explore module interactions and their cumulative effects. Experimental results confirm that each module improves intention prediction, with additional gains from their integration.
Figure 6: F1-scores (%) of various attention mechanisms on the Brain4Cars dataset, offset by 15 units for visual clarity.
Effect of attention design in CaTFormer.
To evaluate the effect of different attention mechanisms on driving intention prediction, we conducted ablation experiments on the Brain4Cars dataset comparing three attention schemes, as summarized in Fig. 6. The results indicate that our proposed Bidirectional Dependency Attention (BDA) more effectively captures spatio-temporal correlations between interior and exterior streams while suppressing noise, thereby demonstrating its superiority in isolating dynamic cues and enhancing overall model robustness.
@@
We study the effect of the intention-loss weight $\alpha$ through an ablation analysis. As shown in Fig. 7, $\alpha=0.10$ achieves the best F1-score, balancing temporal learning with intention supervision and providing informative gradients without destabilizing the main objective. Smaller $\alpha$ under-supervises subtle pre-action cues, whereas larger $\alpha$ induces gradient conflicts that degrade temporal coherence and anticipation accuracy.
@@ Conclusion
In this paper, we introduced CaTFormer, a unified architecture that explicitly models causal interactions between driver behavior and environmental context for accurate intention prediction. Our approach extracts dual-stream features and merges them through a structured fusion pipeline. Extensive experiments on the Brain4Cars dataset confirm that CaTFormer achieves state-of-the-art accuracy, demonstrating its suitability for real-time driver assistance.
Acknowledgments
This work was supported in part by the Beijing Jiaotong University Research Fund under Grant KKA309004533.